# A Dynamic Model of Performative Human-ML Collaboration: Theory and Empirical Evidence

### v2

@@ Abstract
Machine learning (ML) models are increasingly used in various applications, from recommendation systems in e-commerce to diagnosis prediction in healthcare. In this paper, we present a novel dynamic framework for thinking about the deployment of ML models in a performative, human-ML collaborative system. In our framework, the introduction of ML recommendations changes the data generating process of human decisions, which are only a proxy to the ground truth and which are then used to train future versions of the model. We show that this dynamic process in principle can converge to different stable points, i.e. where the ML model and the Human+ML system have the same performance. Some of these stable points are suboptimal with respect to the actual ground truth. We conduct an empirical user study with 1,408 participants to showcase this process. In the study, humans solve instances of the knapsack problem with the help of machine learning predictions. This is an ideal setting because we can see how ML models learn to imitate human decisions and how this learning process converges to a stable point. We find that for many levels of ML performance, humans can improve the ML predictions to dynamically reach an equilibrium performance that is around 92% of the maximum knapsack value. We also find that the equilibrium performance could be even higher if humans rationally followed the ML recommendations. Finally, we test whether monetary incentives can increase the quality of human decisions, but we fail to find any positive effect. Our results have practical implications for the deployment of ML models in contexts where human decisions may deviate from the indisputable ground truth.
@@ 1 Introduction
Human-ML collaboration is increasingly used in various applications, from content moderation in social media [1] to predicting diagnoses in healthcare [2, 3] and making hiring decisions in human resources [4]. Companies that implement human-ML collaborative systems face three crucial challenges: 1) ML models learn from past human decisions, which are often only an approximation to the ground truth (noisy labels); 2) ML models are rolled out to help future human decisions, affecting the data-generating process of human-ML collaboration that then influences future updates to ML models (performative predictions [5]); and 3) the quality of the human-ML collaborative prediction of the ground truth may change as a function of incentives and other human factors. These challenges create a dynamic learning process. Without access to the ground truth, it is often difficult to know whether the learning process will reach an equilibrium state with a good approximation of the ground truth, if it is interrupted at a sub-optimal level, or if it does not reach a stable state at all.
@@
This raises a multitude of empirical questions regarding the governing mechanisms of this dynamic system. How do humans improve on ML predictions of different quality levels, and do financial incentives matter? Will the dynamic learning process converge to a good equilibrium even without the company knowing the actual ground truth labels? How do we design an experiment to test such a system in an empirical context?
@@
In this paper, we present a novel framework for thinking about ML deployment strategies in a performative, human-AI collaborative system. We present a simple theoretical framework to identify conditions under which ML deployment strategies converge to stable points that are a good approximation to the ground truth, and conditions under which there are downward spirals away from the ground truth. To validate our theory, we provide an empirical study in which humans solve knapsack problems with the help of machine learning predictions. We conducted a user study with 1,408 participants, each of whom solved 10 knapsack problems. Knapsack problems are particularly suited to explore our questions because we can train machine learning models to replicate human decisions, while having access to the underlying ground truth (the optimal solution) to evaluate the learning process. Additionally, knapsack problems can be hard for humans, making the task not obvious; and problem instances can be generated and labeled perfectly at negligible cost.
We highlight both theoretical and empirical contributions. Our theoretical framework introduces the collaborative characteristic function as the function mapping the performance of ML models with respect to the ground truth to the performance of humans using those models when making decisions (human+ML). We define utility the performance of decisions (made by ML alone, human alone, or human+ML) with respect to the ground truth. We note that utility cannot be easily quantified for many practical applications. Finally, we introduce the notion of collaborative learning paths, each of which characterizes a possible dynamic deployment strategy. We show conditions under which this dynamic system theoretically reaches a stable point of utility for the firm.
On the empirical side, we show how low-quality ML models can make humans perform worse than if they had no access to ML recommendations. Still, humans improve on ML models so that the deployment strategy of ML models with initial performance between 72% and 92% will converge to a performance that is around 92% of the value of the optimal knapsack solution. Two empirical findings constrain the system from converging to an even higher performance. First, humans do not respond to financial incentives for performance. Second, humans sometimes make decisions that are worse than the ML recommendation, despite the fact that it is fairly easy for them to compare their solution to the ML suggestion and pick the best of the two.
@@ 2 Related Work
There has been a growing body of work investigating various forms of human-AI collaboration. From learning-to-defer systems, where a model defers prediction tasks to humans if its own uncertainty is too high[6, 7, 8], to AI-assisted decision making where humans may or may not consult ML predictions to make a decision [9, 3, 2]. Several alternative decision mechanisms have also been explored [10, 11]. The application areas range from programming [12, 13], to healthcare [2, 3] and business consulting [14]. Related work also investigates factors influencing human-ML collaboration, such as explanations of ML predictions [15], monetary incentives [16], fairness constraints [17], and humans’ adaptability to model changes [18]. In this work, for the first time to the best of our knowledge, we examine the human+ML interaction from a dynamic perspective, where ML models learn from human decisions that are 1) the result of previous human+ML collaboration and 2) can arbitrarily deviate from the underlying ground truth.
This paper is also inspired by an extensive line of work on performative prediction [5, 19, 20, 21], a theoretical framework in which predictions influence the outcome they intend to predict. We adapt the ideas of performative prediction to a context of human-ML collaboration and extend it in three major ways: 1) In our setting, the model predictions change the quality of the human-ML labels as a proxy for the ground truth (e.g., a doctor diagnosis), but the ground truth is held constant (e.g., the true patient diagnosis); 2) We introduce the concept of utility, to quantify the quality of a solution with respect to the ground truth. There can be several stable points with respect to model parameters in the performative prediction framework, but not all of them are equally good at approximating the indisputable ground truth; and 3) The ground truth is unknown, and the mapping between human or ML labels and the ground truth is not fixed. To the best of our knowledge, we are the first to explore performative predictions where the company deploying ML models is unaware of the model’s performance relative to the ground truth, and only knows its similarity to human labels. Our empirical application is also novel in that it investigates the implications of performative predictions for human-ML collaboration.
@@ 3 Problem Statement
We consider a setting in which time is separable in discrete time epochs $t=1,...,T$. At each $t$, a firm deploys machine learning model $M_{t}\in\mathcal{M}$ with $M_{t}:\mathcal{X}\rightarrow\mathcal{Y}$. The model $M_{t}$ predicts a solution $Y\in\mathcal{Y}$ (e.g., a diagnosis) to a problem $X\in\mathcal{X}$ (e.g. the patient’s symptoms) as a function of past data. The firm employs expert humans $H\in\mathcal{H}$ with $H:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{Y}$, who solve the problems with the help of the ML predictions. We will write $M_{t}(X)=Y_{M_{t}}$ and $H(X,Y_{M_{t}})=Y_{H_{t}}$. We assume that for all $X\in\mathcal{X}$, there exists an optimal solution $Y^{*}$, which is the undisputable ground truth.
@@
In many real world applications, determining the ground truth label $Y^{*}$ can be extremely costly. For example, obtaining the correct medical diagnosis can often require the knowledge of various specialists (e.g., orthopedists, pediatricians, neurologists). Even when a single expert is enough, they can misdiagnose a patient’s symptoms. Yet, in many of these cases, using the human labels $Y_{H_{t}}$ as a proxy for $Y^{*}$ is the only feasible option to build ML models. We allow for the quality of $Y_{H_{t}}$ with respect to $Y^{*}$ to change. This means that two iterations of the ML model, $M_{t}$ and $M_{t+1}$, are trained on data from two different data generating processes, $(X,Y_{H_{t-1}})\sim~{}D_{t-1}$ and $(X,Y_{H_{t}})\sim D_{t}$, respectively.
Without access to $Y^{*}$, the only feasible learning objective for a firm that wants to update its model parameters at time $t$ is the comparison between the latest human-ML collaborative labels with the new predictions. [^1] We assume that models at time $t$ are trained exclusively on data from the previous period $t-1$, although we can generalize our setting to include any data points from 0 to $t-1$. For a given loss function $\mathit{l}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R_{+}}$ that is
@@
For simplicity, we assume that at each time $t$, the firm collects enough data to perfectly learn the human-ML solution. In other words, with the optimal model, $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We return to this assumption in Appendix A.7.
@@
As time $t$ increases, the firm hopes that the distributions $D_{t}$ shift closer to the optimal distribution $D^{*}$, where $(X,Y)=(X,Y^{*})$. In other words, for each model’s distance $d$, $d(D_{t},D^{*})>d(D_{t+1},D^{*})$. This could happen, for example, if humans were able to easily compare available solutions and pick the one that is closest to the indisputable ground truth. We can translate this continuous improvement into properties of the human decision function $H$ as follows: for all $t=1,...,T$ and $X\in\mathcal{X}$,
@@
The firm’s hope is that $\delta_{M_{t}}\geq 0$ for $M_{t}$. Effectively, $\delta_{M_{t}}$ characterizes the human-ML collaboration for all utility levels of a model. If $\delta_{M_{t}}$ is positive, humans are able to improve on a ML prediction (and future model iterations will thus get better at approximating the ground truth). Instead, if $\delta_{M_{t}}$ is negative, humans will perform worse than the ML recommendations, and future model iterations will get progressively farther away from the ground truth. We define the function given by Equation 3 as the collaborative characteristic function:
@@
(Collaborative Characteristic Function) For a utility function $\mathbb{U}$, humans $H\in\mathcal{H}$, we define the collaborative characteristic function $\Delta_{\mathbb{U}}:\mathbb{R}\rightarrow\mathbb{R}$ as follows:
$\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))=\underset{H\in\mathcal{H}}{\mathbb{E}}[\mathbb{U}(X,H(X,Y_{M}))]=\mathbb{U}(X,Y_{M})+\delta_{M}.$
The function $\Delta_{\mathbb{U}}$ can take any arbitrary form. Several factors can affect $\Delta_{\mathbb{U}}$, e.g., monetary incentives and ML explanations. In subsequent sections, we empirically approximate one such function.
@@
In what follows, we empirically explore a context where 1) the ML model learns to imitate human decisions, and 2) it is easy for us to identify the indisputable ground truth. The setting allows us to draw at least a portion of the collaborative characteristic function, and explore the effects of human behavior on its shape, particularly the effect of monetary incentives and alternative selection criteria. We present study participants with instances of hard knapsack problems to answer the following research questions:
@@
RQ2: What is the shape of the human-ML collaborative characteristic function $\Delta_{\mathbb{U}}$? Here, we hold the performance bonus constant, and test humans’ effect $\delta_{M}$ on utility for different levels of ML performance. This will enable us to approximate $\Delta_{\mathbb{U}}$ for a specific task.
@@ 4 Experimental Setup
The knapsack problem has desirable properties for the empirical application of our dynamic framework. First, users do not require special training—beyond a short tutorial—to find a solution to the problem. Yet, the task is hard for humans, especially with a growing number of items [22]. Thus, the optimal solution $Y^{*}$ is not obvious. Second, we can generate solutions to the knapsack problem in two ways. The “optimal” solution can be found with dynamic programming. The “ML” solution can be found by imitating what humans select and computing the training loss as the difference between the items selected by participants versus items selected by a model.
@@
Appendix A.4 contains details about $\mathbb{U}_{\text{Econ}}(X,Y)$ and discusses our results using an alternative utility function. Note that there can be multiple optimal combinations of items to pack, but the optimal value $Y^{*}$ is unique.
Study Design.
@@
We ran the 10-cent Bonus treatment twice. In the second round, we introduced a comprehension quiz to ensure that our participants understood the payment structure. Within the 10-cent Bonus with bonus comprehension quiz, we randomized access to ML recommendations. Users were randomly allocated to one of seven ML treatments. The control group had no ML recommendations. The other six groups had access to recommendations from a progressively better ML model, as Table 1 shows on each of the last six columns. We picked ML models with varying degrees of performance to approximate the collaborative characteristic function from Figure 1. For details on the model training for the ML recommendations, see Appendix A.8.

### v3

@@ Abstract
Machine learning (ML) models are increasingly used in various applications, from recommendation systems in e-commerce to diagnosis prediction in healthcare. In this paper, we present a novel dynamic framework for thinking about the deployment of ML models in a performative, human-ML collaborative system. In our framework, the introduction of ML recommendations changes the data-generating process of human decisions, which are only a proxy to the ground truth and which are then used to train future versions of the model. We show that this dynamic process in principle can converge to different stable points, i.e. where the ML model and the Human+ML system have the same performance. Some of these stable points are suboptimal with respect to the actual ground truth. As a proof of concept, we conduct an empirical user study with 1,408 participants. In the study, humans solve instances of the knapsack problem with the help of machine learning predictions of varying performance. This is an ideal setting because we can identify the actual ground truth, and evaluate the performance of human decisions supported by ML recommendations. We find that for many levels of ML performance, humans can improve upon the ML predictions. We also find that the improvement could be even higher if humans rationally followed the ML recommendations. Finally, we test whether monetary incentives can increase the quality of human decisions, but we fail to find any positive effect. Using our empirical data to approximate our collaborative system suggests that the learning process would dynamically reach an equilibrium performance that is around 92% of the maximum knapsack value. Our results have practical implications for the deployment of ML models in contexts where human decisions may deviate from the indisputable ground truth.
@@ 1 Introduction
Human-ML collaboration is increasingly used in various applications, from content moderation in social media (Lai et al., 2022) to predicting diagnoses in healthcare (Jacobs et al., 2021; Dvijotham et al., 2023) and making hiring decisions in human resources (Peng et al., 2022). Companies that implement human-ML collaborative systems face three crucial challenges: 1) ML models learn from past human decisions, which are often only an approximation to the ground truth (noisy labels); 2) ML models are rolled out to help future human decisions, affecting the data-generating process of human-ML collaboration that then influences future updates to ML models (performative predictions as in Perdomo et al. (2020)); and 3) the quality of the human-ML collaborative prediction of the ground truth may change as a function of incentives and other human factors. These challenges create a dynamic learning process. Without access to the ground truth, it is often difficult to know whether the learning process will reach an equilibrium state with a good approximation of the ground truth, if it is interrupted at a sub-optimal level, or if it does not reach a stable state at all.
@@
This raises a multitude of empirical questions regarding the governing mechanisms of this dynamic system. How do humans improve on ML predictions of different quality levels, and do financial incentives matter? Will the dynamic learning process converge to a good equilibrium even without the company knowing the actual ground truth labels?
@@
In this paper, we present a novel framework for thinking about ML deployment strategies in a performative, human-AI collaborative system. We present a theoretical framework to identify conditions under which ML deployment strategies converge to stable points that are a good approximation to the ground truth, and conditions under which there are downward spirals away from the ground truth. Our theory introduces the notion of a collaborative characteristic function, which maps algorithmic performance to the performance of human decisions supported by ML predictions. As a proof of concept for our theory, we provide an empirical study in which humans solve knapsack problems with the help of machine learning predictions. We conducted a user study with 1,408 participants, each of whom solved 10 knapsack problems. The empirical exercise allows us to evaluate 1) the quality of human decisions supported by ML models of varying performance, 2) how these human decisions compare to a best-case scenario, and 3) how these human decisions are affected by monetary incentives. With some additional assumptions, we can map these data to our theoretical collaborative characteristic function.
We highlight three main empirical results. First, we show that humans tend to improve upon the ML recommendation for many levels of ML performance. Second, humans sometimes submit solutions that are worse than the ML recommendation, despite the fact that with knapsack, it is fairly easy for them to compare their solution to the ML suggestion and pick the best of the two. Third, humans do not respond to financial incentives for performance. The empirical data can be used to approximate theoretical collaborative characteristic functions. The results suggest that, at least in our context, collaborative characteristic functions are invariant to monetary incentives. Additionally, the fact that humans sometimes submit solutions that are worse than the provided ML recommendation implies that there remains a gap between the collaborative characteristic function based on human labels and the collaborative characteristic function constructed by selecting the maximum between the human and ML solution.
@@ 2 Related Work
There has been a growing body of work investigating various forms of human-ML collaboration. From learning-to-defer systems, where a model defers prediction tasks to humans if its own uncertainty is too high (Cortes et al., 2016; Charusaie et al., 2022; Mozannar et al., 2023), to ML-assisted decision making where humans may or may not consult ML predictions to make a decision (Mozannar et al., 2024c; Dvijotham et al., 2023; Jacobs et al., 2021). Several alternative decision mechanisms have also been explored (Steyvers et al., 2022; Mozannar et al., 2024a). The application areas range from programming (Dakhel et al., 2023; Mozannar et al., 2024b), to healthcare (Jacobs et al., 2021; Dvijotham et al., 2023) and business consulting (Dell’Acqua et al., 2023). Related work also investigates factors influencing human-ML collaboration, such as explanations of ML predictions (Vasconcelos et al., 2023), monetary incentives (Agarwal et al., 2023), fairness constraints (Sühr et al., 2021), and humans’ adaptability to model changes (Bansal et al., 2019). In this work, for the first time to the best of our knowledge, we theoretically examine the human+ML interaction from a dynamic perspective, where ML models learn from human decisions that are 1) the result of previous human+ML collaboration and 2) can arbitrarily deviate from the underlying ground truth.
This paper is also inspired by an extensive line of work on performative prediction (Perdomo et al., 2020; Mendler-Dünner et al., 2020; Hardt et al., 2022; Mendler-Dünner et al., 2022), a theoretical framework in which predictions influence the outcome they intend to predict. We adapt the ideas of performative prediction to a context of human-ML collaboration and extend it in three major ways: 1) In our setting, the model predictions change the quality of the human-ML labels as a proxy for the ground truth (e.g., a doctor diagnosis), but the ground truth is held constant (e.g., the true patient diagnosis); 2) We introduce the concept of utility, to quantify the quality of a solution with respect to the ground truth. There can be several stable points with respect to model parameters in the performative prediction framework, but not all of them have the same utility, i.e., are equally good at approximating the indisputable ground truth; 3) The ground truth is unknown, and the mapping between human or ML labels and the ground truth is not fixed. To the best of our knowledge, we are the first to explore performative predictions where the deployment of ML models occurs while the model’s performance relative to the ground truth is unknown, and only its similarity to human labels is available. Our empirical application is also novel in that it provides a first step towards investigating the implications of performative predictions for human-ML collaboration.
@@ 3 Problem Statement
We consider a setting in which time is separable in discrete time epochs $t=1,...,T$. At each $t$, a firm deploys machine learning model $M_{t}\in\mathcal{M}$ of a model class $\mathcal{M}$, with $M_{t}:\mathcal{X}\rightarrow\mathcal{Y}$. The model $M_{t}$ predicts a solution $Y\in\mathcal{Y}$ (e.g., a diagnosis) to a problem $X\in\mathcal{X}$ (e.g. the patient’s symptoms) as a function of past data. The firm employs expert humans $H\in\mathcal{H}$ with $H:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{Y}$, who solve the problems with the help of ML predictions. We will write $M_{t}(X)=Y_{M_{t}}$ and $H(X,Y_{M_{t}})=Y_{H_{t}}$. We assume that for all $X\in\mathcal{X}$, there exists an optimal solution $Y^{*}$, which is the indisputable ground truth.
@@
In many real-world applications, determining the ground truth label $Y^{*}$ can be extremely costly. For example, obtaining the correct medical diagnosis can often require the knowledge of various specialists (e.g., orthopedists, pediatricians, neurologists). Even when a single expert is enough, they can misdiagnose a patient’s symptoms. Yet, in many of these cases, using the human labels $Y_{H_{t}}$ as a proxy for $Y^{*}$ is the only feasible option to build ML models. We allow the quality of $Y_{H_{t}}$ with respect to $Y^{*}$ to change. This means that two iterations of the ML model, $M_{t}$ and $M_{t+1}$, are trained on data from two different data generating processes, $(X,Y_{H_{t-1}})\sim~{}D_{t-1}$ and $(X,Y_{H_{t}})\sim D_{t}$, respectively.
Without access to $Y^{*}$, the only feasible learning objective for a firm that wants to update its model parameters at time $t$ is the comparison between the latest human-ML collaborative labels with the new predictions. [^1] We assume that models at time $t$ are trained exclusively on data from the previous period $t-1$, although we can generalize our setting to include any data points from 0 to $t-1$. For a given loss function $\mathit{l}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R_{+}}$ we can write this as follows:
@@
For simplicity, we assume that at each time $t$, the firm collects enough data to perfectly learn the human-ML solution. In other words, with the optimal model, $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We discuss relaxing this assumption in Appendix A.7.
@@
As time $t$ increases, the firm hopes that the distributions $D_{t}$ shift closer to the optimal distribution $D^{*}$, where $(X,Y)=(X,Y^{*})$. In other words, for each model’s distance $d$, $d(D_{t},D^{*})>d(D_{t+1},D^{*})$. This could happen, for example, if humans were able to easily compare available solutions and pick the one that is closest to the indisputable ground truth.
We can translate this continuous improvement into properties of the human decision function $H$ as follows: for all $t=1,...,T$ and $X\in\mathcal{X}$,
@@
The firm’s hope is that $\delta_{M_{t}}\geq 0$ for $M_{t}$. Effectively, $\delta_{M_{t}}$ characterizes the human-ML collaboration for all utility levels of a model. If $\delta_{M_{t}}$ is positive, humans are able to improve on a ML prediction (and future model iterations will thus get better at approximating the ground truth). Instead, if $\delta_{M_{t}}$ is negative, humans will perform worse than the ML recommendations, and future model iterations will get progressively farther away from the ground truth.
We define the function given by Equation 3 as the collaborative characteristic function:
@@
(Collaborative Characteristic Function) For a utility function $\mathbb{U}$, humans $H\in\mathcal{H}$ and model $M\in\mathcal{M}$, we define the collaborative characteristic function $\Delta_{\mathbb{U}}:\mathbb{R}\times\mathcal{M}\rightarrow\mathbb{R}$ as follows:
$\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}),M)=\underset{H\in\mathcal{H}}{\mathbb{E}}[\mathbb{U}(X,H(X,Y_{M}))]=\mathbb{U}(X,Y_{M})+\delta_{M}.$
The function $\Delta_{\mathbb{U}}$ can take any arbitrary form. Several factors can affect $\Delta_{\mathbb{U}}$, e.g., ML explanations and monetary incentives (as we empirically explore in Section 4). Note that $\Delta_{\mathbb{U}}$ is also a function of the ML model $M$, and not just of its utility, because equal levels of utility across different ML models do not guarantee equal collaborative reactions from humans. In the rest of the paper however, we will shorten the notation and write $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))$.
@@
In what follows, we offer a proof of concept of our theoretical setup. We empirically explore a context where it is easy for us to identify the indisputable ground truth. Although with some simplifying assumptions, the setting allows us to approximate a portion of the collaborative characteristic function, and explore the effects of human behavior on its shape, particularly the effect of monetary incentives and alternative solution selection criteria. We present study participants with instances of hard knapsack problems to answer the following research questions:
@@
RQ2: Can we approximate the human-ML collaborative characteristic function $\Delta_{\mathbb{U}}$? Here, we hold the performance bonus constant, and test humans’ effect $\delta_{M}$ on utility for different levels of ML performance. This will enable us to construct two approximations of $\Delta_{\mathbb{U}}$ for a specific task.
@@ 4 Experimental Setup
The knapsack problem has desirable properties for the empirical application of our framework. First, users do not require special training—beyond a short tutorial—to find a solution to the problem. Yet, the task is hard for humans, especially with a growing number of items (Murawski & Bossaerts, 2016). Thus, the optimal solution $Y^{*}$ is not obvious. Second, we can generate solutions to the knapsack problem in two ways. The “optimal” solution can be found with dynamic programming. The “ML” solution can be found by imitating what humans select and computing the training loss as the difference between the items selected by participants versus items selected by a model. Finally, it allows us to showcase our theoretical approach with two different utility functions and two selection criteria to approximate the collaborative characteristic function.
@@
Appendix A.4 contains details about $\mathbb{U}_{\text{Econ}}(X,Y)$ and discusses our results using an alternative utility function $\mathbb{U}_{\text{Opt}}(X,Y)$ (optimality), which is equal to one if a solution is optimal and zero otherwise. Note that there can be multiple optimal combinations of items to pack, but the optimal value $Y^{*}$ is always unique.
@@
Study Design.
A total of 1,408 participants completed the study; we removed 119 participants due to forbidden browser reloads or uses of the browser’s back-button, which left 1,289 for the analyses below. See Appendix A.9 for an overview of participants’ demographics. On average, participants’ compensation implied an hourly wage of £12.17 (\$15.22), which is above the UK minimum wage of £11.44. Appendix A.3 contains additional payment details.
@@
We ran the 10-cent Bonus treatment twice. In the second round, we introduced a comprehension quiz to ensure that our participants understood the payment structure. Within the 10-cent Bonus with bonus comprehension quiz, we randomized access to ML recommendations. Users were randomly allocated to one of seven ML treatments. The control group had no ML recommendations. The other six groups had access to recommendations from progressively better ML models, denoted q1 through q6 as Table 1 shows on each of the last six columns.
The rationale for selecting the treatment conditions described above is the following. First, we want to understand whether monetary incentives change human effort, which in turn would translate into a shift in the collaborative characteristic function from Figure 1. Although ideally one would want to approximate the entire collaborative characteristic function under different incentive structures, to ensure statistical power under a limited budget, we opted for testing the role of varying bonuses without ML recommendations. Second, we want to understand how ML models of varying performance affect the human-ML performance to draw the collaborative characteristic function. Ideally, one would want these models to be trained on human labels (themselves potentially affected by previous model iterations) to mirror the theoretical framework, but this would have required sequential rounds of experimentation. Instead, we approximate the sequential nature of our framework by training all models on optimally solved knapsack instances, rather than instances solved by humans. Appendix A.8 discusses further details on the model training. Ex-post, we verify that models trained on human labels (from the data collected during the experiment) have a wide range of utility levels. Appendix Figure 6 confirms that the utilities of the models selected for our treatments fall comfortably within the large range of utility levels of models trained on human labels. Nonetheless, we emphasize that these design choices limit our ability to truly replicate our theoretical model. We return to the limitations of this approach in the conclusion.